# Negative Dependence as a toolbox for machine learning : review and new developments

### v1

@@ 4.2.2 Monte Carlo methods
Monte Carlo methods [RC05] refers to picking up the nodes as the realization of random points in $\mathbb{R}^{d}$. A standard algorithm in this vein is the so-called importance sampling. We consider the case when $\mu(\mathrm{d}x)=\omega(x)\mathrm{d}x$ supported in the hypercube $I^{d}$. Importance sampling outputs $N$ nodes $X_{1},\ldots,X_{N}$ as i.i.d. samples from a proposal density $q(x)\mathrm{d}x$ and
@@ 4.3 Monte Carlo with DPPs
In [BH20], the authors show that using DPPs for Monte Carlo methods can give better decay rates for $\mathrm{Var}[\mathcal{E}_{N}(f)]$. Their approach is based on the theory of (multivariate) orthogonal polynomials.
@@
has mean zero by construction. Intuitively, the negative dependence in DPPs results in more cancellation in linear statistics, often leading to smaller variance. Motivated by this intuition, the authors in [BH20] demonstrated that this is indeed the case for the multivariate OPE through a careful analysis of the estimator (4.1). To convey their idea while maintaining simplicity, we will present a particular case from their results.
Theorem 4.1 (Theorem 1 in [BH20]).
@@
In fact, the authors further demonstrated that the rate mentioned above eventually applies to a larger class of measures $\mu$, and presented central limit theorems suited for Monte Carlo integration. For a more general result, we refer readers to the paper [BH20].
@@ 5.1 A parametric model
Directionality in data pertains to isotropy-breaking – namely, the degree of dependency between points (for instance, repulsion) is much stronger in a few special directions compared to others. Thus, directionality corresponds to a type of (low dimensional) structure on data (compare, e.g., to sparsity), and is naturally applicable to learning scenarios such as clustering where such distinguished directions are of canonical importance. In [GR20] the authors considered Gaussian Determinantal Processes (abbrv. GDP) as a well-structured parametric model of DPPs, with particular focus on exploring isotropy versus directionality in data.
@@
$K(x-y)=\frac{1}{(2\pi)^{\frac{d}{2}}\sqrt{\mathrm{Det}{\Sigma}}}\exp\big(-\frac{1}{2}(x-y)^{\top}\Sigma^{-1}(x-y)\big{missing})\,,\quad x,y\in\mathbb{R}^{d}\,.$
@@ 5.2 A spiked model for DPPs and directionality in data
$\bar{\rho}_{2}(x,y)=-\exp\big(-(x-y)^{T}\Sigma^{-1}(x-y)\big{missing})\,.$
Therefore, if $x-y$ is well-aligned with an eigenvector of $\Sigma$ corresponding to a relatively large eigenvalue, then the magnitude of $\bar{\rho}_{2}(x,y)$ remains significant even when $\|x-y\|$ is large. As such, in these directions, the spatial correlation extends over a much longer range (compared to others).
This observation forms the basis of a spiked DPP model, analogous to the widely-used spiked covariance model [BBAP05], where the covariance matrix $\Sigma$ is modeled as a rank-one perturbation of the identity: $\Sigma=\text{Id}+\lambda uu^{\top}$, with $\|u\|=1$ and $\lambda\geq 0$. In this GDP setting, the directional structure is captured by the rank-one perturbation. In the context of GDPs, it is natural to consider scenarios where the presence of the spike does not alter the mean density of points compared to the isotropic case, where the scattering matrix is the identity. If this condition does not hold, the spike’s presence can be detected by estimating the density. Maintaining the same mean density implies that $\mathrm{Det}(\Sigma)$ is the same for both the null and alternative cases; this corresponds to shear-type transformations of the ambient Euclidean space.
Putting together the above observations, [GR20] proposes the following spiked model for DPPs:
@@ 5.3 Inference on spiked models for GDP
Inferential procedures on the GDP, as laid out in [GR20], are based on statistic $\hat{\Sigma}$ of given by the $d\times d$ matrix
@@
and denoting by $\chi(B)$ the indicator of the event $B$, it is demonstrated in [GR20] that the test
@@
detects the presence of a spike with power $\delta$ as soon as the spike strength $\lambda$ is above the threshold $\lambda_{\mathrm{thresh}}=2\mathfrak{r}_{n,d}$. In the event that a spike is detected, the spike direction can be effectively estimated by the maximal eigenvector of $\hat{\Sigma}$. It has been demonstrated in [GR20] that these statistical procedures achieve parametric rates in fixed dimensions. This raises natural and interesting questions vis-a-vis the study of GDP type models and related statistical procedures in a high dimensional setting, wherein new kinds of random matrix ensembles are expected to emerge.
@@ 5.4 Clustering and random matrices
It is well known that the principal directions in PCA correspond to the directions of maximal variance in the data. In [GR20], dimension reduction methodology based on finding directions of maximal repulsion was proposed, based on the GDP as an ansatz. In this vein, we obtain dimension reduction by projecting onto the principal eigendirections of the matrix $\hat{\Sigma}$ as above. The performance of this approach has been demonstrated on the celebrated Fisher’s Iris dataset, which is a prototypical benchmark for clustered data [GR20].
Figure 2: Left: Iris-DPP     Right: W-DPP   (c.f. [GR20])
@@
It maybe noted that, even if the data points $\{X_{1},\ldots,X_{n}\}$ are independent, random matrices of the form $\tilde{\Sigma}$ have dependent entries (due to the truncation in the definition). However, as a random matrix model, it still offers a significant degree of structure so that hope for analytical progress is still reasonable. Indeed, in [GMT24], the authors investigate the most basic setting for such kernelized random matrix ensembles, where the points $\{X_{1},\ldots,X_{n}\}$ are i.i.d., and establish the emergence of novel families limiting spectral distributions that generalize the classical Marcenko-Pastur laws, and beyond. The methods involve crucial ingredients from free probability theory. We expect that, motivated by the GDP ansatz and its methodological application, the study of such kernelized ensembles will lead to novel and interesting problems in random matrix theory with dependent entries.
@@ 6.1 Coresets in machine learning
In machine learning, learning tasks are often formulated as an optimization problem of minimizing a suitable loss function [SSBD14]. Let $\mathcal{X}$ be a set of size $N$ (called the data set) and $\mathcal{F}$ be a family of real-valued functions on $\mathcal{X}$. In this survey, we will focus on a loss function of the following form
@@
=$\mathcal{F}=\{(y,z)\mapsto|\langle a,y\rangle+b-z|^{2},a\in\mathbb{R}^{d},b\in\mathbb{R}\}.$
When the data size $N$ is large, the computational cost can become prohibitive, and the optimization problem often becomes intractable. A natural approach is to reduce the amount of data by using only a manageable number of representative samples. This idea is formalised by the notion of coreset (see [BLK17]).
@@ 6.2 DPPs for minibatch sampling in SGD
where $\mathcal{X}:=\{x_{1},\ldots,x_{N}\}$ is the data set, and the loss $\ell_{\theta}(\cdot)$ is differentiable in the parameter $\theta$. When $N\gg 1$, the computational cost for the gradient of the objective function becomes expensive. SGD method [RM51] refers to building an estimator for the gradient at each iteration of the gradient descent, using a small samples of data points (called a minibatch). Given $m\ll N$, a key problem in SGD is to construct minibatches of size $m$ that keep the variance of the gradient estimators as small as possible [MB11, ZKM17]. More precisely, we want to construct a (random) subset $\mathcal{S}\subset\mathcal{X}$ of size $m$, such that
@@
In [BGL21], the authors used DPPs to construct minibatch sampling for SGD which yields significant variance reduction, compared to independent sampling. To elaborate, we assume that the data $x_{i}$’s are drawn i.i.d. from a distribution $\gamma(x)\mathrm{d}x$ on the hypercube $[-1,1]^{d}$. Using a suitable multivariate OPE and spectral approximation, Bardenet et al. constructed an $N\times N$ orthogonal projection matrix $\mathbf{K}$ of rank $m$. Let $\mathcal{S}$ be the DPP on $\mathcal{X}$ defined by the kernel matrix $\mathbf{K}$ w.r.t. the reference measure $N^{-1}\sum_{i=1}^{N}\delta_{x_{i}}$, the estimator is then given by
@@
For more detailed discussions, we refer the readers to their paper [BGL21] and the references therein. We remark that, for independent sampling minibatches of size $m$, the variance of $L_{\mathcal{S}}(\theta)$ is of order $m^{-1}$. Thus, sampling with DPPs improves the decay rate of the variance in the exponent, which is significant.
@@ 6.3 Importance sampling for coresets
We come back to the problem of randomized coresets construction. A well-studied method to construct coresets is the so-called importance sampling ([BLK17], Chapter 2). Given a proposal probability distribution $q$ on $\mathcal{X}$ and $m\in\mathbb{N}$, importance sampling outputs $\mathcal{S}_{\mathrm{iid}}=\{X_{1},\ldots,X_{m}\}$, where $X_{i}\sim_{\mathrm{iid}}q$, and weights $w(x)=(mq(x))^{-1}$. Thus,
@@
A natural task is then to find a distribution $q$ that makes $\mathrm{Var}[L_{\mathcal{S}_{\mathrm{iid}}}(f)]$ small uniformly in $f\in\mathcal{F}$. A candidate for this task is the so-called sensitivity based importance sampling, introduced by Langberg and Schulman [LS10]. For simplicity, let us consider the case when $\mathcal{F}$ consists of non-negative functions.
@@
$\mathbb{P}\Big{(}\Big{|}\frac{L_{\mathcal{S}_{\mathrm{iid}}}(f)}{L(f)}-1\Big{|}>\varepsilon\Big{)}\leq 2\exp\Big(-\frac{2m\varepsilon^{2}}{S}\Big{missing})\cdot$
The sharper the bound $s(x)\geq\sigma(x)$, the better performance of importance sampling. Obtaining a sharp upper bound on sensitivity could be a challenge and will depend on the specific models (see [BLK17]).
After establishing the concentration result for each $L_{\mathcal{S}}(f)$, the next step is to obtain a uniform bound over $\mathcal{F}$. For this task, it is necessary to quantify the complexity of $\mathcal{F}$, and the uniform bound will depend heavily on that quantity. In [BLK17], the authors use Vapnik-Chervonenkis dimension [VC71] and pseudo-dimension to quantify the complexity of $\mathcal{F}$, and then apply a chaining argument to obtain the desired upper bound. However, it is important to note that this chaining argument relies on the i.i.d. nature of the sampling scheme. For more details, we refer the readers to [BLK17].
@@ 6.4 DPPs as coresets
Intuitively, DPPs avoid situations where a data point is sampled multiple times, leading to redundancy. Furthermore, as we have already observed, DPPs with suitable kernels often result in smaller variance for linear statistics ([BH20, BGL21]). Thus, there is hope to improve performance by using DPPs instead of independent sampling.
In [TBA19], Tremblay et al. provide extensive theoretical and empirical justification for the use of DPPs to construct randomized coresets. However, the crucial question of whether DPP-based coresets can provide a strict improvement remained open. As mentioned earlier, constructing coresets necessitates concentration inequalities and chaining, and these theories have been well-developed for independent processes. Unfortunately, corresponding results for correlated processes are less established, which could pose an obstacle for sampling with correlated processes.
In a recent result [BGST24], the authors confirm that, with a careful choice of kernels, DPP-based coresets can outperform independently drawn coresets. The key ingredient is a Bernstein-type concentration inequality for linear statistics of DPPs. This result was first investigated in a seminal paper [BD14] by Breuer and Duits for projection DPPs. In [BGST24], the authors extend this result to the case of DPPs with general Hermitian kernels, and even non-symmetric kernels, which seems to go beyond the state of the art. It is worth noting that DPPs with non-symmetric kernels have recently been shown to be of significant interest in machine learning, but come with a limited theoretical toolbox [GBDK19, GHD+20, HGDK22].
@@ 6.4.1 Concentration inequalities for linear statistics of DPPs
In [PP14], Pemantle and Peres introduced concentration inequalities for Lipschitz functionals of probability measures on $2^{\mathcal{X}}$ (equivalently, random subsets of $\mathcal{X}$) satisfying the stochastic covering property (SCP). We refer the readers to [PP14] to the precise definition of SCP. However, we remark that the class of probability measures on $2^{\mathcal{X}}$ satisfying the SCP particularly includes DPPs.
For simplicity, we only state here the result for $m$-homogeneous point processes (i.e., point processes with a.s. $m$ points). For the general case, we refer the readers to [PP14].
Theorem 6.7 (Theorem 3.1 in [PP14]).
@@
$\mathbb{P}(F(\mathcal{S})-\mathbb{E}[F(\mathcal{S})]\geq a)\leq\exp\Big(-\frac{a^{2}}{8m}\Big{missing}),\quad\forall a\geq 0.$
@@
$\mathbb{P}(\Lambda_{\mathcal{S}}(f)-\mathbb{E}[\Lambda_{\mathcal{S}}(f)]\geq a)\leq\exp\Big(-\frac{a^{2}}{8m\|f\|_{\infty}^{2}}\Big{missing}),\quad\forall a\geq 0.$
@@
for some constant $C>0$ (possibly depending on $f$), which is not better than independent sampling (c.f. Proposition 6.6). One possible reason is that the concentration inequalities in Theorem 6.7 and Proposition 6.8 do not take the variance into account; meanwhile the variance reduction for linear statistics of DPPs is one of the main motivation for sampling with DPPs. Thus, a concentration inequality for linear statistics of DPPs which takes variance into account is a desire. As we introduced earlier, such results have been investigated in [BD14, BGST24].
@@
$\mathbb{P}(|\Lambda_{\mathcal{S}}(f)-\mathbb{E}\Lambda_{\mathcal{S}}(f)|\geq a)\leq 2\exp\Big(-\frac{a^{2}}{4A\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]}\Big{missing}),\quad\forall 0\leq a\leq\frac{2A\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]}{3\|f\|_{\infty}},$
@@
$\mathbb{P}(|\Lambda_{\mathcal{S}}(f)-\mathbb{E}[\Lambda_{\mathcal{S}}(f)]|\geq a)\leq 2\exp\Big(-{a^{2}\over 4\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]}\Big{missing}),\>\forall 0\leq a\leq{\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]^{2}\over 40\|f\|_{\infty}^{3}\cdot\max(1,\|K\|_{\mathrm{op}}^{2})\cdot\|K\|_{*}},$
@@
$\mathbb{P}\Big{(}\Big{|}\frac{L_{\mathcal{S}}(f)}{L(f)}-1\Big{|}\geq\varepsilon\Big{)}\leq 2\exp\Big(-\frac{\varepsilon^{2}}{4A}\mathrm{Var}\Big{[}\frac{L_{\mathcal{S}}(f)}{L(f)}\Big{]}^{-1}\Big{missing}),$
@@
To compare, assume that $\mathbb{E}[|\mathcal{S}|]=m$, then typically $\mathrm{Var}[L_{\mathcal{S}}(f)/L(f)]=O(m^{-(1+\delta)})$, for some parameter $\delta>0$, depending on the specific model (see [BGL21], [BH20]). Hence, Corollary 6.12 yields a concentration inequality of the form

### v4

@@ 4.2.2 Monte Carlo methods
Monte Carlo methods [MonteCarlobook] refers to picking up the nodes as the realization of random points in $\mathbb{R}^{d}$. A standard algorithm in this vein is the so-called importance sampling. We consider the case when $\mu(\mathrm{d}x)=\omega(x)\mathrm{d}x$ supported in the hypercube $I^{d}$. Importance sampling outputs $N$ nodes $X_{1},\ldots,X_{N}$ as i.i.d. samples from a proposal density $q(x)\mathrm{d}x$ and
@@ 4.3 Monte Carlo with DPPs
In [OPE-AOAP], the authors show that using DPPs for Monte Carlo methods can give better decay rates for $\mathrm{Var}[\mathcal{E}_{N}(f)]$. Their approach is based on the theory of (multivariate) orthogonal polynomials.
@@
has mean zero by construction. Intuitively, the negative dependence in DPPs results in more cancellation in linear statistics, often leading to smaller variance. Motivated by this intuition, the authors in [OPE-AOAP] demonstrated that this is indeed the case for the multivariate OPE through a careful analysis of the estimator (4.1). To convey their idea while maintaining simplicity, we will present a particular case from their results.
Theorem 4.1 (Theorem 1 in [OPE-AOAP]).
@@
In fact, the authors further demonstrated that the rate mentioned above eventually applies to a larger class of measures $\mu$, and presented central limit theorems suited for Monte Carlo integration. For a more general result, we refer readers to the paper [OPE-AOAP].
@@ 5.1 A parametric model
Directionality in data pertains to isotropy-breaking – namely, the degree of dependency between points (for instance, repulsion) is much stronger in a few special directions compared to others. Thus, directionality corresponds to a type of (low dimensional) structure on data (compare, e.g., to sparsity), and is naturally applicable to learning scenarios such as clustering where such distinguished directions are of canonical importance. In [GDP] the authors considered Gaussian Determinantal Processes (abbrv. GDP) as a well-structured parametric model of DPPs, with particular focus on exploring isotropy versus directionality in data.
@@
$K(x-y)=\frac{1}{(2\pi)^{\frac{d}{2}}\sqrt{\mathrm{Det}{\Sigma}}}\exp\big(-\frac{1}{2}(x-y)^{\top}\Sigma^{-1}(x-y)\big.)\,,\quad x,y\in\mathbb{R}^{d}\,.$
@@ 5.2 A spiked model for DPPs and directionality in data
[…]
Therefore, if xyx-y is well-aligned with an eigenvector of $\Sigma$ corresponding to a relatively large eigenvalue, then the magnitude of $\bar{\rho}_{2}(x,y)$ remains significant even when $\|x-y\|$ is large. As such, in these directions, the spatial correlation extends over a much longer range (compared to others).
This observation forms the basis of a spiked DPP model, analogous to the widely-used spiked covariance model [BBP], where the covariance matrix $\Sigma$ is modeled as a rank-one perturbation of the identity: $\Sigma=\text{Id}+\lambda uu^{\top}$, with $\|u\|=1$ and $\lambda\geq 0$. In this GDP setting, the directional structure is captured by the rank-one perturbation. In the context of GDPs, it is natural to consider scenarios where the presence of the spike does not alter the mean density of points compared to the isotropic case, where the scattering matrix is the identity. If this condition does not hold, the spike’s presence can be detected by estimating the density. Maintaining the same mean density implies that $\mathrm{Det}(\Sigma)$ is the same for both the null and alternative cases; this corresponds to shear-type transformations of the ambient Euclidean space.
Putting together the above observations, [GDP] proposes the following spiked model for DPPs:
@@ 5.3 Inference on spiked models for GDP
Inferential procedures on the GDP, as laid out in [GDP], are based on statistic $\hat{\Sigma}$ of given by the $d\times d$ matrix
@@
and denoting by $\chi(B)$ the indicator of the event $B$, it is demonstrated in [GDP] that the test
@@
detects the presence of a spike with power $\delta$ as soon as the spike strength $\lambda$ is above the threshold $\lambda_{\mathrm{thresh}}=2\mathfrak{r}_{n,d}$. In the event that a spike is detected, the spike direction can be effectively estimated by the maximal eigenvector of $\hat{\Sigma}$. It has been demonstrated in [GDP] that these statistical procedures achieve parametric rates in fixed dimensions. This raises natural and interesting questions vis-a-vis the study of GDP type models and related statistical procedures in a high dimensional setting, wherein new kinds of random matrix ensembles are expected to emerge.
@@ 5.4 Clustering and random matrices
It is well known that the principal directions in PCA correspond to the directions of maximal variance in the data. In [GDP], dimension reduction methodology based on finding directions of maximal repulsion was proposed, based on the GDP as an ansatz. In this vein, we obtain dimension reduction by projecting onto the principal eigendirections of the matrix $\hat{\Sigma}$ as above. The performance of this approach has been demonstrated on the celebrated Fisher’s Iris dataset, which is a prototypical benchmark for clustered data [GDP].
Figure 2: Left: Iris-DPP     Right: W-DPP   (c.f. [GDP])
@@
It maybe noted that, even if the data points $\{X_{1},\ldots,X_{n}\}$ are independent, random matrices of the form $\tilde{\Sigma}$ have dependent entries (due to the truncation in the definition). However, as a random matrix model, it still offers a significant degree of structure so that hope for analytical progress is still reasonable. Indeed, in [ghosh-mukherjee-talukdar], the authors investigate the most basic setting for such kernelized random matrix ensembles, where the points $\{X_{1},\ldots,X_{n}\}$ are i.i.d., and establish the emergence of novel families limiting spectral distributions that generalize the classical Marcenko-Pastur laws, and beyond. The methods involve crucial ingredients from free probability theory. We expect that, motivated by the GDP ansatz and its methodological application, the study of such kernelized ensembles will lead to novel and interesting problems in random matrix theory with dependent entries.
@@ 6.1 Coresets in machine learning
In machine learning, learning tasks are often formulated as an optimization problem of minimizing a suitable loss function [Shalev-Shwartz_Ben-David_2014]. Let $\mathcal{X}$ be a set of size $N$ (called the data set) and $\mathcal{F}$ be a family of real-valued functions on $\mathcal{X}$. In this survey, we will focus on a loss function of the following form
@@
$\mathcal{F}=\{(y,z)\mapsto|\langle a,y\rangle+b-z|^{2},a\in\mathbb{R}^{d},b\in\mathbb{R}\}.$
When the data size $N$ is large, the computational cost can become prohibitive, and the optimization problem often becomes intractable. A natural approach is to reduce the amount of data by using only a manageable number of representative samples. This idea is formalised by the notion of coreset (see [bachem2017coresetML]).
@@ 6.2 DPPs for minibatch sampling in SGD
where $\mathcal{X}:=\{x_{1},\ldots,x_{N}\}$ is the data set, and the loss $\ell_{\theta}(\cdot)$ is differentiable in the parameter $\theta$. When $N\gg 1$, the computational cost for the gradient of the objective function becomes expensive. SGD method [Robbins_SGD] refers to building an estimator for the gradient at each iteration of the gradient descent, using a small samples of data points (called a minibatch). Given $m\ll N$, a key problem in SGD is to construct minibatches of size $m$ that keep the variance of the gradient estimators as small as possible [NIPS2011_SGD, zhang2017_SGD]. More precisely, we want to construct a (random) subset $\mathcal{S}\subset\mathcal{X}$ of size $m$, such that
@@
In [OPE-NIPS], the authors used DPPs to construct minibatch sampling for SGD which yields significant variance reduction, compared to independent sampling. To elaborate, we assume that the data $x_{i}$’s are drawn i.i.d. from a distribution $\gamma(x)\mathrm{d}x$ on the hypercube $1,1]^{d}$. Using a suitable multivariate OPE and spectral approximation, Bardenet et al. constructed an $N\times N$ orthogonal projection matrix $\mathbf{K}$ of rank $m$. Let $\mathcal{S}$ be the DPP on $\mathcal{X}$ defined by the kernel matrix $\mathbf{K}$ w.r.t. the reference measure $N^{-1}\sum_{i=1}^{N}\delta_{x_{i}}$, the estimator is then given by
@@
For more detailed discussions, we refer the readers to their paper [OPE-NIPS] and the references therein. We remark that, for independent sampling minibatches of size $m$, the variance of $L_{\mathcal{S}}(\theta)$ is of order $m^{-1}$. Thus, sampling with DPPs improves the decay rate of the variance in the exponent, which is significant.
@@ 6.3 Importance sampling for coresets
We come back to the problem of randomized coresets construction. A well-studied method to construct coresets is the so-called importance sampling ([bachem2017coresetML], Chapter 2). Given a proposal probability distribution $q$ on $\mathcal{X}$ and $m\in\mathbb{N}$, importance sampling outputs $\mathcal{S}_{\mathrm{iid}}=\{X_{1},\ldots,X_{m}\}$, where $X_{i}\sim_{\mathrm{iid}}q$, and weights $w(x)=(mq(x))^{-1}$. Thus,
@@
A natural task is then to find a distribution $q$ that makes $\mathrm{Var}[L_{\mathcal{S}_{\mathrm{iid}}}(f)]$ small uniformly in $f\in\mathcal{F}$. A candidate for this task is the so-called sensitivity based importance sampling, introduced by Langberg and Schulman [sensitivity]. For simplicity, let us consider the case when $\mathcal{F}$ consists of non-negative functions.
@@
[…]
The sharper the bound $s(x)\geq\sigma(x)$, the better performance of importance sampling. Obtaining a sharp upper bound on sensitivity could be a challenge and will depend on the specific models (see [bachem2017coresetML]).
After establishing the concentration result for each $L_{\mathcal{S}}(f)$, the next step is to obtain a uniform bound over $\mathcal{F}$. For this task, it is necessary to quantify the complexity of $\mathcal{F}$, and the uniform bound will depend heavily on that quantity. In [bachem2017coresetML], the authors use Vapnik-Chervonenkis dimension [VCdim] and pseudo-dimension to quantify the complexity of $\mathcal{F}$, and then apply a chaining argument to obtain the desired upper bound. However, it is important to note that this chaining argument relies on the i.i.d. nature of the sampling scheme. For more details, we refer the readers to [bachem2017coresetML].
@@ 6.4 DPPs as coresets
Intuitively, DPPs avoid situations where a data point is sampled multiple times, leading to redundancy. Furthermore, as we have already observed, DPPs with suitable kernels often result in smaller variance for linear statistics ([OPE-AOAP, OPE-NIPS]). Thus, there is hope to improve performance by using DPPs instead of independent sampling.
In [tremblay2019determinantal], Tremblay et al. provide extensive theoretical and empirical justification for the use of DPPs to construct randomized coresets. However, the crucial question of whether DPP-based coresets can provide a strict improvement remained open. As mentioned earlier, constructing coresets necessitates concentration inequalities and chaining, and these theories have been well-developed for independent processes. Unfortunately, corresponding results for correlated processes are less established, which could pose an obstacle for sampling with correlated processes.
In a recent result [NIPS2024], the authors confirm that, with a careful choice of kernels, DPP-based coresets can outperform independently drawn coresets. The key ingredient is a Bernstein-type concentration inequality for linear statistics of DPPs. This result was first investigated in a seminal paper [BREUER2014441] by Breuer and Duits for projection DPPs. In [NIPS2024], the authors extend this result to the case of DPPs with general Hermitian kernels, and even non-symmetric kernels, which seems to go beyond the state of the art. It is worth noting that DPPs with non-symmetric kernels have recently been shown to be of significant interest in machine learning, but come with a limited theoretical toolbox [gartrell2019learning, gartrell2020scalable, han2022scalable].
@@ 6.4.1 Concentration inequalities for linear statistics of DPPs
In [PEMANTLE_PERES_2014], Pemantle and Peres introduced concentration inequalities for Lipschitz functionals of probability measures on $2^{\mathcal{X}}$ (equivalently, random subsets of $\mathcal{X}$) satisfying the stochastic covering property (SCP). We refer the readers to [PEMANTLE_PERES_2014] to the precise definition of SCP. However, we remark that the class of probability measures on $2^{\mathcal{X}}$ satisfying the SCP particularly includes DPPs.
For simplicity, we only state here the result for $m$-homogeneous point processes (i.e., point processes with a.s. $m$ points). For the general case, we refer the readers to [PEMANTLE_PERES_2014].
Theorem 6.7 (Theorem 3.1 in [PEMANTLE_PERES_2014]).
@@
$\mathbb{P}(F(\mathcal{S})-\mathbb{E}[F(\mathcal{S})]\geq a)\leq\exp\Big(-\frac{a^{2}}{8m}\Big.),\quad\forall a\geq 0.$
@@
$\mathbb{P}(\Lambda_{\mathcal{S}}(f)-\mathbb{E}[\Lambda_{\mathcal{S}}(f)]\geq a)\leq\exp\Big(-\frac{a^{2}}{8m\|f\|_{\infty}^{2}}\Big.),\quad\forall a\geq 0.$
@@
for some constant $C>0$ (possibly depending on $f$), which is not better than independent sampling (c.f. Proposition 6.6). One possible reason is that the concentration inequalities in Theorem 6.7 and Proposition 6.8 do not take the variance into account; meanwhile the variance reduction for linear statistics of DPPs is one of the main motivation for sampling with DPPs. Thus, a concentration inequality for linear statistics of DPPs which takes variance into account is a desire. As we introduced earlier, such results have been investigated in [BREUER2014441, NIPS2024].
@@
$\mathbb{P}(|\Lambda_{\mathcal{S}}(f)-\mathbb{E}\Lambda_{\mathcal{S}}(f)|\geq a)\leq 2\exp\Big(-\frac{a^{2}}{4A\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]}\Big.),\quad\forall 0\leq a\leq\frac{2A\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]}{3\|f\|_{\infty}},$
@@
$\mathbb{P}(|\Lambda_{\mathcal{S}}(f)-\mathbb{E}[\Lambda_{\mathcal{S}}(f)]|\geq a)\leq 2\exp\Big(-{a^{2}\over 4\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]}\Big.),\>\forall 0\leq a\leq{\mathrm{Var}[\Lambda_{\mathcal{S}}(f)]^{2}\over 40\|f\|_{\infty}^{3}\cdot\max(1,\|K\|_{\mathrm{op}}^{2})\cdot\|K\|_{*}},$
@@
[…]
@@
To compare, assume that $\mathbb{E}[|\mathcal{S}|]=m$, then typically $\mathrm{Var}[L_{\mathcal{S}}(f)/L(f)]=O(m^{-(1+\delta)})$, for some parameter $\delta>0$, depending on the specific model (see [OPE-NIPS], [OPE-AOAP]). Hence, Corollary 6.12 yields a concentration inequality of the form